# GINTRIP: Interpretable Temporal Graph Regression using Information bottleneck and Prototype-based method

### v1

@@ Abstract
Deep neural networks (DNNs) have demonstrated remarkable performance across various domains, yet their application to temporal graph regression tasks faces significant challenges regarding interpretability. This critical issue, rooted in the inherent complexity of both DNNs and underlying spatio-temporal patterns in the graph, calls for innovative solutions. While interpretability concerns in Graph Neural Networks (GNNs) mirror those of DNNs, to the best of our knowledge, no notable work has addressed the interpretability of temporal GNNs using a combination of Information Bottleneck (IB) principles and prototype-based methods. Our research introduces a novel approach that uniquely integrates these techniques to enhance the interpretability of temporal graph regression models. The key contributions of our work are threefold:
We introduce the Graph INterpretability in Temporal Regression task using Information bottleneck and Prototype (GINTRIP) framework, the first combined application of IB and prototype-based methods for interpretable temporal graph tasks. We derive a novel theoretical bound on mutual information (MI), extending the applicability of IB principles to graph regression tasks. We incorporate an unsupervised auxiliary classification head, fostering multi-task learning and diverse concept representation, which enhances the model bottleneck’s interpretability.
Our model is evaluated on real-world traffic datasets, outperforming existing methods in both forecasting accuracy and interpretability-related metrics. The implementation of our research is available at Github.
@@
00footnotetext: Equal contribution
Figure 1: Proposed method architecture. It includes an encoder, learnable prototypical layer, subgraph extractor, regression, and classification head components.
@@ I Introduction
In the intelligent transportation industry, ensemble tree-based methods such as XGBoost [1] are in high demand. Their popularity is due to their interpretability. Meanwhile, with the success of Graph Neural Networks (GNNs) in a wide range of deep learning tasks, there has been an increasing demand to explore the decision-making process of these models and provide explanations for their predictions. Existing traffic dashboards such as STRADA[2] cannot interpret the predictions. To meet this requirement, explainable AI (XAI) has developed as a means to interpret black-box models by offering clear explanations for their outputs.
Interpretability in deep neural networks (DNNs) is one of the most important neglected aspects that still need more investigation due to the black-box nature of DNNs. Different tools and methods have been used to address this black box. One of them is the IB [3] that is raised when analyzing DNNs using information theoretic concepts. On the other hand, some other techniques, like prototype-based methods, use learnable blocks that can capture key features that are useful for interpretability[4, 5]. For instance, [6] attempted to optimize IB bounds via adversarial training for semantic segmentation tasks while the interpretability is yet neglected. On the other hand, [7] tackles the same task with the notion of prototypes, leading to explainability.
In the area of graphs, there are prominent works to integrate IB. Specifically, [8, 9] are the pioneers who derive tractable bounds to achieve interpretable Graph Neural Networks (GNNs). In the terminology of interpretability, they are classified into the post-hoc methods, which help to understand and interpret the predictions of black-box models like deep neural networks [10],[11], which are often opaque in terms of decision-making processes. On the other hand, another class of interpretable models are self-explainable methods. One such approach is to utilize the IB principle to explore the natural interpretability and generalization capabilities of GNNs [12]. Furthermore, recent progress in intrinsic approaches [13] has tackled explainability issues and addressed challenges in managing graph out-of-distribution cases by applying invariant learning and causal inference techniques. In the temporal graph domain, IB methods are unexplored. One of the main works in this area is [14], which generates spatial and temporal subgraphs as explainability output. These subgraphs are not linked to prototypes, hindering high-level interpretability.
@@
  • We are the first model to perform interpretable temporal graph regression based on the prototype notion
  • we provide pseudo-labels to facilitate interpretability
  • Our framework derives tractable MI bounds for the temporal graph regression problem.
  • Our model is extensively evaluated on real-world traffic datasets, for which it outperforms the state-of-the-art.
@@ II-A Problem Statement
We focus on temporal graph regression, a technique for predicting time-dependent node values in graph-structured data. We apply this to traffic forecasting, which anticipates traffic conditions based on historical data. In our graph model $G_{in}=(V_{G_{in}},E,A)$. Here, $V_{G_{in}}=\left\{V_{1},V_{2},...,V_{N}\right\}$ is the node set, $E$ the edge set, and $A\in\mathbb{R}^{N\times N}$ the adjacency matrix. At each timestamp $t$, the graph $G_{in}$ includes a dynamic feature matrix $X_{t}\in\mathbb{R}^{N\times D}$. In this study, the graph is static and the traffic feature is speed. Given a length-${T^{\prime}}$ forecast horizon and a length-$T$ observed series history $\left[y_{1},...,y_{T}\right]\in\mathbb{R}^{T}$, the goal is to predict the vector of future values. For simplicity, we consider a sliding window of length $W\leq T$ ending with the most recent observed value $y_{T}$ as the model input, denoted as $\boldsymbol{x}\in\mathbb{R}^{W}=\left[y_{T-W+1},y_{T-W+2},...,y_{T}\right]$. The prototype layer is also incorporated into our model. It is defined as a set ${G}_{p}=\left\{\mathbf{z}_{G_{p}}^{1},\mathbf{z}_{G_{p}}^{2},\cdots,\mathbf{z}_{G_{p}}^{M}\right\}$, where $M$ is the total number of prototypes, and each prototype $\mathbf{z}_{G_{p}}^{m}$ is a learnable parameter vector that acts as the latent representation of the prototypical part (i.e., ${G}_{p}$) of graph ${G_{in}}$. We allocate $J$ prototypes for each pseudo-class, i.e., $M=K\times J$, where $K$ denotes the number of classes corresponding to our unsupervised pseudo-labels. Indeed, for each node $V_{i}$, a threshold $T_{i}=Q_{0.1}(X_{:,i,:})$ classifies each node into one of our predefined pseudo-classes(congested/non-congested).
@@ II-C Prototype Injection
Prototype-based methods are used in different machine-learning tasks to capture the essential key of the bottleneck variable of the input. Combining this technique with GIB could provide and extract a more meaningful subgraph of the input graph. Ultimately, we develop the prototype layer as a learnable component to discern predictive temporal patterns. To do so, we compute the similarity of the extracted subgraph and the prototype matrix as a predictive feature during forecasting. Using $G_{p}$, we can inject $G_{p}$ trainable variables to the first term of (3) i.e. $I(Y;G_{sub})$. Based on the chain rule of MI, we can write: $I(Y;G_{sub})=I(Y;G_{sub},G_{p})-I(Y;G_{p}|G_{sub})$ , and then we can rewrite (3): $\small\min_{G_{sub}}-I(Y;G_{sub},G_{p})+I(Y;G_{p}|G_{sub})+\beta I(G_{in};G_{sub})$. Due to some difficulty in computing the second term, we again use the chain rule of MI. We can decompose the second term, and we have a more extended expression:
@@
Observing the first two terms in (4), we obviously recognize that they are opposite, meaning the first MI wants to increase, and the second one wants to decrease. But, after enough training time, these two terms force each other, where an increase in $I(Y;G_{sub},G_{p})$ leads to an increase in the $I(G_{p};Y,G_{sub})$ that is different behavior than (4). So, we exclude the second MI term from our optimization objective to prevent this misalignment:
@@
After sufficient training, $I(G_{p};G_{sub})$ increases, indicating stronger correlation between $G_{p}$ and $G_{sub}$. Fig. 1 shows the relationship complexity of $G_{p}$ and $G_{sub}$ is lower than that of $G_{sub}$ and $Y=(Y_{cls},Y_{reg})$. Assuming $G_{p}\approx G_{sub}$, we get $I(G_{p};G_{sub})\approx H(G_{sub})$, leading to $I(G_{p};G_{sub})\gtrsim I(Y;G_{sub})$. Consequently, $I(G_{p};Y,G_{sub})\gtrsim I(Y;G_{sub},G_{p})$, implying an increase in $I(Y;G_{sub},G_{p})$ leads to an undesirable increase in $I(G_{p};Y,G_{sub})$, conflicting with optimization (4).
@@ II-D Subgraph Extraction
Our model adopted Spatio-Temporal Graph Neural Networks (STGNN) to process the input temporal graph. More specifically, we used MTGNN[15] as our GNN backbone. Following [16], we passed the nodes’ extracted features, referred to as $h_{i}$ in Fig. 1, to the sequence of an MLP layer and a sigmoid activation function. Their output assigns the probability $p_{i}$ of the node inclusion in the selected subgraph, output of the $G_{sub}$-Extractor block in Fig. 1. The final representation of each selected subgraph node would be: $\small z_{i}=\lambda_{i}h_{i}+(1-\lambda_{i})\epsilon$, where $\lambda_{i}\sim Bernoulli(p_{i})$ and $\epsilon\sim\mathcal{N}(\mu_{h_{i}},\sigma_{h_{i}}^{2})$. The learned probability $p_{i}$ facilitates selective information retention in $G_{sub}$. This approach preserves interpretability within the subgraph while also potentially aiding the learning of prototypes introduced in the previous step. To minimize the MI between the input graph and the selected subgraph, [17] showed that it is equivalent to minimizing the following upper bound for $I(G_{in},G_{sub})$:
$\mathbb{E}_{G_{in}}\left[-\frac{1}{2}\log S+\frac{1}{2|V_{G_{in}}|}S+\frac{1}{2|V_{G_{in}}|}M^{2}\right]\triangleq\mathcal{L}_{sub}$ (6)
@@
To enforce learning connected subgraphs[16], we employ the following loss function, which is called connectivity loss: $\small\mathcal{L}_{con}\triangleq\left\|\widehat{P^{T}AP}-I_{2}\right\|_{F}$. where $P\in\mathcal{R}^{\left|V_{G}\right|\times 2}$ and $A$ are the node probability assignment and adjacency matrix at the batch level, respectively. $I_{2}$ is the 2x2 identity matrix, $\|...\|_{F}$ is the Frobenius norm operation and $\widehat{(...)}$ is the row normalization. More precisely, each $P$ matrix row is related to the probability of its corresponding node, meaning that for each row we have ($p_{i},1-p_{i}$). Notably, locality preservation is a crucial aspect of the extracted subgraph[19].
[…]
@@ II-E Theoretical Bounds
To optimize our objective, we must consider the inherent characteristics of each MI term. By applying appropriate lower and upper bounds to these terms, we define a constrained optimization problem that forms our final loss function. Below, we derive lower bounds for the first two terms of the above optimization objective. $I(Y_{cls};G_{sub},G_{p})$ is equal:
$\mathbb{E}_{Y_{cls},G_{sub},G_{p}}[\log p(Y_{cls}|G_{sub},G_{p})]-\mathbb{E}_{Y_{cls}}[\log p(Y_{cls})]$ (7)
@@
Where $\gamma$ is a similarity function between $G_{sub},G_{p}$. The first inequality was derived from DPI (Data Process Inequality)[20] and the second one, caused by $q_{\phi}(Y_{cls}|\gamma(G_{sub},G_{p}))$, which is the variational approximation to the true posterior $p(Y_{cls}|\gamma(G_{sub},G_{p}))$ Equation (7) illustrates that maximizing the MI $I(Y_{cls};G_{sub},G_{p})$ is equivalent to minimizing the classification loss, $\mathcal{L}_{cls}$. In practice, the cross-entropy loss is typically employed for categorical $Y_{cls}$.
@@
Where $\lambda$ is a constant coefficient and also $\hat{Y}_{reg}$ and $Y_{reg}$ are the classification head prediction and the classification pseudo-label (ground-truth), respectively (refer to Fig. 1).
Following a similar variational approach to that used in equation (8), we can express:
$I(G_{sub};G_{p})\geq-\lambda^{\prime}MSE(Z_{sub},Z_{G_{p}})\triangleq-\lambda^{\prime}\mathcal{L}_{var}$ (9)
Here, $\lambda^{\prime}$ is a constant coefficient. $Z_{sub}$ denotes the output of the pooling function applied to $G_{sub}$, while $Z_{G_{p}}$ represents the concatenation of all prototype vectors $\left\{\mathbf{z}_{{G}_{p}}^{1},\mathbf{z}_{{G}_{p}}^{2},\cdots,\mathbf{z}_{{G}_{p}}^{M}\right\}$. For $q_{\phi}$, we employ a single-layer perceptron to estimate $Z_{G_{p}}$ from $Z_{sub}$. The total loss function is a convex combination of individual losses derived from the aforementioned bounds: $\mathcal{L}=\lambda_{1}\mathcal{L}_{reg}+\lambda_{2}\mathcal{L}_{sub}+\lambda_{3}\mathcal{L}_{var}+\lambda_{4}\mathcal{L}_{con}+\lambda_{5}\mathcal{L}_{cls}$. To address the instability of this function during training, we employed the multi-loss variation coefficient adjustment approach [21].

### v2

@@ Abstract
Deep neural networks (DNNs) have demonstrated remarkable performance across various domains, but their inherent complexity makes them challenging to interpret. This is especially true for temporal graph regression tasks due to the complex underlying spatio-temporal patterns in the graph.
While interpretability concerns in Graph Neural Networks (GNNs) mirror those of DNNs, no notable work has addressed the interpretability of temporal GNNs to the best of our knowledge. Innovative methods, such as prototypes, aim to make DNN models more interpretable. However, a combined approach based on prototype-based methods and Information Bottleneck (IB) principles has not yet been developed for temporal GNNs. Our research introduces a novel approach that uniquely integrates these techniques to enhance the interpretability of temporal graph regression models. The key contributions of our work are threefold:
We introduce the Graph INterpretability in Temporal Regression task using Information bottleneck and Prototype (GINTRIP) framework, the first combined application of IB and prototype-based methods for interpretable temporal graph tasks. We derive a novel theoretical bound on mutual information (MI), extending the applicability of IB principles to graph regression tasks. We incorporate an unsupervised auxiliary classification head, fostering diverse concept representation using multi-task learning, which enhances the model’s interpretability.
Our model is evaluated on real-world datasets like traffic and crime, outperforming existing methods in both forecasting accuracy and interpretability-related metrics such as MAE, RMSE, MAPE, and fidelity. The code is available on Github.
@@
[…]
=-21pt
@@ I Introduction
The rapid expansion of intelligent vehicles, driven by socioeconomic factors and urbanization, underscores the critical necessity for Intelligent Transportation Systems to mitigate the resulting challenges of escalating urban congestion and environmental degradation [1][2][3][4][5].
In the intelligent transportation industry, ensemble tree-based methods such as XGBoost [6] are popular for the regression task thanks to their interpretability. Meanwhile, with the success of Graph Neural Networks (GNNs) in a wide range of deep learning tasks, there has been an increasing demand to explore the decision-making process of these models and provide explanations for their predictions. Existing traffic dashboards such as STRADA [7] cannot interpret their predictions.
Interpretability is one of the most important yet often neglected aspects of deep neural networks (DNNs) that still need more investigation due to their black-box nature. Different tools and methods have been used to unveil this black box. One of them is the Information Bottleneck (IB) [8], a principle for analyzing DNNs using information theoretic concepts. The main advantage of the IB is that a DNN model filters out the noise and extracts informative latent representations that are easier to understand and analyze. This compression facilitates the identification of which aspects of the input data are most influential in the model’s predictions, thereby enhancing interpretability[8]. On the other hand, techniques like prototype-based methods use learnable blocks that can capture key features that are useful for interpretability [9, 10]. For instance, [11] attempted to optimize IB bounds via adversarial training for semantic segmentation tasks while the interpretability was neglected. On the other hand, [12] tackled the same task with the notion of prototypes, leading to interpretability.
In the area of graphs, there are prominent works to integrate IB. Specifically, [13, 14] are the pioneers who derive tractable bounds to achieve interpretable GNNs. In the terminology of interpretability, one can distinguish the post-hoc methods, which help to understand and interpret the predictions of black-box models like deep neural networks [15, 16], which are often opaque in terms of decision-making processes. Another class of interpretable models is that of the self-explainable methods. One such approach is to utilize the IB principle to explore the natural interpretability and generalization capabilities of GNNs [17]. Furthermore, recent progress in intrinsic approaches [18] has tackled explainability issues and addressed challenges in managing graph out-of-distribution cases by applying invariant learning and causal inference techniques. In the temporal graph domain, IB methods are underexplored. One of the main works in this area is [19], which generates spatial and temporal subgraphs as explainability output. These subgraphs are not linked to prototypes, hindering high-level interpretability.
@@
  • We are the first model to perform interpretable temporal graph regression based on the prototype notion.
  • We provide pseudo-labels to facilitate interpretability.
  • Our framework derives tractable mutual information (MI) bounds for the temporal graph regression problems.
  • Our model is extensively evaluated on real-world traffic and crime datasets, for which it outperforms the state-of-the-art.
Figure 1: Proposed method architecture. It includes a GNNGNN encoder, subgraph extractor, learnable prototypes, regression, and classification head components. Overall, the temporal graph passes to the GNNGNN model, and the latent representation $h_{1},h_{2},...,h_{N}$} is processed by $G_{sub}$ ExtractorExtractor to extract a subgraph $G_{sub}$. A pooling block is used to capture global-level features. A similarity function $\gamma$ computes similarities between learnable prototypes ${G}_{p}=\left\{\mathbf{z}_{G_{p}}^{1},\mathbf{z}_{G_{p}}^{2},\cdots,\mathbf{z}_{G_{p}}^{M}\right\}$ and subgraph representation $Z_{sub}$. The concatenated similarities $S^{1},S^{2},...,S^{M}$}} with $Z_{sub}$ are used for the proposed pseudo-label classification and the forecasting task.
@@ II-A Problem Statement
We focus on temporal graph regression, a technique for predicting time-dependent node values in graph-structured data. We apply this to the datasets, which anticipate traffic and crime conditions based on historical data. In our graph model $G_{in}=(V_{G_{in}},E_{G_{in}},A_{G_{in}})$. Here, $V_{G_{in}}=\left\{V_{1},V_{2},...,V_{N}\right\}$ is the set of nodes, $E_{G_{in}}$ the set of edges, and $A_{G_{in}}\in\mathbb{R}^{N\times N}$ the adjacency matrix. At each timestamp $t$, the graph $G_{in}$ includes a dynamic feature matrix $X_{t}\in\mathbb{R}^{N\times D}$. In this study, the graph structure is static, and the node’s features are attributes related to each dataset, e.g., traffic speed in the traffic dataset. Given a forecast horizon of length $T^{\prime}$ and an observed series history of length $T$, $\left[y_{1},...,y_{T}\right]\in\mathbb{R}^{T}$, the goal is to predict the vector of future values. For simplicity, we consider a sliding window of length $W\leq T$ ending with the most recent observed value $y_{T}$ as the model input, denoted as $\bm{x}\in\mathbb{R}^{W}=\left[y_{T-W+1},y_{T-W+2},...,y_{T}\right]$, where $\bm{Y}$ is the prediction series. The prototype is defined as a set ${G}_{p}=\left\{\mathbf{z}_{G_{p}}^{1},\mathbf{z}_{G_{p}}^{2},\cdots,\mathbf{z}_{G_{p}}^{M}\right\}$, where $M$ is the total number of prototypes, and each prototype $\mathbf{z}_{G_{p}}^{m}$ is a learnable parameter vector that acts as the latent representation of the prototypical part (i.e., ${G}_{p}$) of graph $G_{in}$}. We allocate $J$ prototypes for each class, i.e., $M=K\times J$, where $K$ denotes the number of classes corresponding to our unsupervised pseudo-classes in the traffic dataset. Indeed, for each node $V_{i}$, a threshold $T_{i}=Q_{0.1}(X_{:,i,:})$ classifies each node into one of our predefined pseudo-classes(congested/non-congested) in the traffic dataset, where $Q(.)$ is quantile function. For the crime dataset, we can use available labels.
@@ II-C Prototype Injection
usepackage[normalem]ulem Prototype-based methods are machine learning techniques that represent data by identifying typical examples, or prototypes, which encapsulate the most relevant features of a dataset. These methods facilitate interpretability by allowing models to make predictions based on the similarity of new inputs to these representative prototypes. In our approach, we integrate the GIB with prototype-based techniques to extract meaningful subgraphs from input graphs. Specifically, we develop a learnable prototype layer designed to identify predictive temporal patterns. This is achieved by computing the similarity between the extracted subgraphs and a prototype matrix, which serves as a predictive feature during the forecasting process. Using $G_{p}$, we can inject $G_{p}$ learnable variables to the first term of (3) i.e. $I(Y;G_{sub})$. Based on the chain rule of MI, we can write:
$\small I(Y;G_{sub})=I(Y;G_{sub},G_{p})-I(Y;G_{p}|G_{sub})$ (4)
and then we can rewrite (3):
$\small\min_{G_{sub}}-I(Y;G_{sub},G_{p})+I(Y;G_{p}|G_{sub})+\beta I(G_{in};G_{sub})$ (5)
Due to some difficulty in computing the second term, we again use the chain rule of MI. We can decompose the second term, and we have a more extended expression:
@@
Observing the first two terms in (6), we recognize that they are opposite, meaning the first MI tends to increase, and the second one tends to decrease. However, after enough training time, these two terms compensate each other, where an increase in $I(Y;G_{sub},G_{p})$ leads to an increase in the $I(G_{p};Y,G_{sub})$ that is different behavior than (6). We therefore exclude the second MI term from our optimization objective to prevent this misalignment:
@@
For an approximate derivation to support the exclusion of this term in (6), at first, we can write:
$\small I(G_{p};Y,G_{sub})=I(Y;G_{sub},G_{p})+I(G_{p};G_{sub})-I(Y;G_{sub})$ (8)
Based on (6), after sufficient training time, $I(G_{p};G_{sub})$ will increase, which means that $G_{p}$ and $G_{sub}$ will have more anafter sufficient trainind enough correlation and dependency. Furthermore, from figure 1, we can see that the relationship complexity of $G_{p}$ and $G_{sub}$ is significantly lower than that of $G_{sub}$ and $Y=(Y_{cls},Y_{reg})$. So for an asymptotic analysis, we can suppose that $G_{p}\approx G_{sub}$, which gives $I(G_{p};G_{sub})\approx H(G_{sub})$ which $H(G_{p})$ is the entropy of $G_{p}$ and then we will have: $I(G_{p};G_{sub})\gtrsim I(Y;G_{sub})$. So, it follows immediately that $I(G_{p};Y,G_{sub})\gtrsim I(Y;G_{sub},G_{p})$ which means an increase in $I(Y;G_{sub},G_{p})$ leads to an increase in $I(G_{p};Y,G_{sub})$. This is an undesirable and conflicting behavior for (6) which intend to decrease; $I(G_{p};Y,G_{sub})$. So, we exclude $I(G_{p};Y,G_{sub})$ from our optimization objective to prevent this misalignment.
@@ II-D Subgraph Extraction
Our model adopts Spatio-Temporal Graph Neural Networks (STGNN) to process the input temporal graph. More specifically, we use MTGNN[20] as our GNN backbone. Following [21], we pass the nodes’ extracted features, referred to as $h_{i}$ in figure 1, to the sequence of an MLP layer and a sigmoid activation function. Their output assigns the probability $p_{i}$ of the node inclusion in the selected subgraph, the output of the $G_{sub}$-Extractor block in figure 1. The final representation of each selected subgraph node would be:
$\small\small z_{i}=\lambda_{i}h_{i}+(1-\lambda_{i})\epsilon$ (9)
, where $\lambda_{i}\sim Bernoulli(p_{i})$ and $\epsilon\sim\mathcal{N}(\mu_{h_{i}},\sigma_{h_{i}}^{2})$. The learned probability $p_{i}$ facilitates selective information retention in $G_{sub}$. This approach preserves interpretability within the subgraph while also potentially aiding the learning of prototypes introduced in the previous step. To minimize the MI between the input graph and the selected subgraph, [22] showed that it is equivalent to minimizing the following upper bound for $I(G_{in},G_{sub})$:
$I(G_{in},G_{sub})\leq\mathbb{E}_{G_{in}}\left[-\frac{1}{2}\log S+\frac{1}{2|V_{G_{in}}|}S+\frac{1}{2|V_{G_{in}}|}M^{2}\right]\triangleq\mathcal{L}_{sub}$ (10)
@@
To enforce learning connected subgraphs [21], we employ the following loss function, which is called connectivity loss:
$\small\mathcal{L}_{con}\triangleq\left\|\widehat{P^{T}AP}-I_{2}\right\|_{F}$ (11)
$P\in\mathcal{R}^{\left|V_{G}\right|\times 2}$ and $A$ are the node probability assignment and adjacency matrix at the batch level, respectively. $I_{2}$ is the 2x2 identity matrix, $\|...\|_{F}$ is the Frobenius norm operation and $\widehat{(...)}$ is the row normalization. More precisely, each $P$ matrix row is related to the probability of its corresponding node, meaning that for each row we have ($p_{i},1-p_{i}$). This preserves locality, which is a crucial aspect of the extracted subgraph [24].
The first MI term in optimization (7), contains the variables’ group $Y=(Y_{cls},Y_{reg})$ which corresponds to our network’s two output heads: classification and regression. Using chain rule, we can derive a desired lower bound: $I(Y;G_{sub},G_{p})\geq\frac{1}{2}I(Y_{cls};G_{sub},G_{p})+\frac{1}{2}I(Y_{reg};G_{sub},G_{p})$. Then, we can rewrite the optimization (7) as:
$\min_{G_{sub}}-\frac{1}{2}I(Y_{cls};G_{sub},G_{p})-\frac{1}{2}I(Y_{reg};G_{sub},G_{p})-I(G_{p};G_{sub})$ (12)
$+\beta I(G_{in};G_{sub})$
@@ II-E Theoretical Bounds
To optimize our objective, we must consider the inherent characteristics of each MI term. By applying appropriate lower and upper bounds to these terms, we define a constrained optimization problem that forms our final loss function 16. Below, we derive lower bounds for the first two terms of the optimization objective 12.
$I(Y_{cls};G_{sub},G_{p})=$ (13)
$\mathbb{E}_{Y_{cls},G_{sub},G_{p}}[\log p(Y_{cls}|G_{sub},G_{p})]-\mathbb{E}_{Y_{cls}}[\log p(Y_{cls})]$
@@
Where $\gamma$ is a similarity function, such as dot product similarity, between $G_{sub},G_{p}$. The first inequality was derived from Data Process Inequality (DPI)[25] and the second one, caused by $q_{\phi}(Y_{cls}|\gamma(G_{sub},G_{p}))$, which is the variational approximation to the true posterior $p(Y_{cls}|\gamma(G_{sub},G_{p}))$ . Equation (13) illustrates that maximizing the MI $I(Y_{cls};G_{sub},G_{p})$ is equivalent to minimizing the classification loss, $\mathcal{L}_{cls}$. In practice, the cross-entropy loss is typically employed for categorical $Y_{cls}$.
@@
Where $\lambda$ is a constant coefficient and also $\hat{Y}_{reg}$ and $Y_{reg}$ are the regression head prediction and the regression actual label (ground-truth), respectively (refer to figure 1).
Following a similar approach to that used in equation (14), we can express:
$I(G_{sub};G_{p})\geq-\lambda^{\prime}MSE(q_{\theta}(Z_{sub}),Z_{G_{p}})\triangleq-\lambda^{\prime}\mathcal{L}_{var}$ (15)
Here, $\lambda^{\prime}$ is a constant coefficient. $Z_{sub}$ denotes the output of the pooling function applied to $G_{sub}$, while $Z_{G_{p}}$ represents the concatenation of all prototype vectors $\left\{\mathbf{z}_{{G}_{p}}^{1},\mathbf{z}_{{G}_{p}}^{2},\cdots,\mathbf{z}_{{G}_{p}}^{M}\right\}$. For $q_{\theta}$, we employ a single-layer perceptron to estimate $Z_{G_{p}}$ from $Z_{sub}$.
The total loss function is a convex combination of individual losses derived from the aforementioned bounds:
$\mathcal{L}=\lambda_{1}\mathcal{L}_{reg}+\lambda_{2}\mathcal{L}_{sub}+\lambda_{3}\mathcal{L}_{var}+\lambda_{4}\mathcal{L}_{con}+\lambda_{5}\mathcal{L}_{cls}$ (16)
To address the instability of this function during training, we employed the multi-loss variation coefficient adjustment approach [26]. Based on this method, each coefficient $\lambda_{s}$ is computed as $\frac{1}{z}\frac{\sigma_{\mathcal{L}_{s}}}{\mu_{\mathcal{L}_{s}}}$, where $\mu_{\mathcal{L}_{s}}$ and $\sigma_{\mathcal{L}_{s}}$ denote the mean and standard deviation of the loss term $\mathcal{L}_{s}$, respectively. The normalizing constant is defined as $z=\sum_{s=1}^{5}\frac{\sigma_{\mathcal{L}_{s}}}{\mu_{\mathcal{L}_{s}}}$, which guarantees $\sum_{s=1}^{5}\lambda_{s}=1$. This constraint is important for decoupling the loss weighting from the learning rate.